# Seoul 121-742, South Korea

### v1

@@ Abstract
Extensions of the Standard Model typically contain “flaton fields” defined as fields with large vacuum expectation values and almost flat potentials where scalar self-coupling is small or vanishes at tree level. Such potentials have been used to drive a secondary inflationary epoch after a primary phase of inflation, in what are called thermal inflation models. Although the primordial, high-scale inflationary epoch can solve the horizon and flatness problems, it does not always resolve difficulties associated with late-time relics produced in extensions of the Standard Model. These relics typically decay too late, injecting entropy and energetic particles that spoil successful predictions like Big Bang Nucleosynthesis. It is here that thermal inflation plays a crucial role: diluting unwanted relics by many orders of magnitude without erasing the baryon asymmetry or the large-scale structure set up by the earlier phase of inflation. The preferred scale for this phenomenon is in the range $10^{6}-10^{8}$ GeV if one considers supergravity, but without it, any scale above the EW scale is valid. We investigate a typical form of these potentials and determine what are the conditions for the potentials to develop a barrier such that when the flatons settle to the true minimum, the associated Gravitational Waves can be observed.
@@ 1 Introduction
Among the most theoretically compelling settings for such a transition is the spontaneous breaking of lepton number symmetry, which dynamically generates Majorana masses for right-handed neutrinos. The observation of neutrino oscillations Fukuda:1998mi ; Abe:2013tm has firmly established that neutrinos have nonzero masses and mix Bilenky:1998dt ; GonzalezGarcia:2002dz ; deSalas:2018bym ; deSalas:2020pgw , a fact that cannot be accommodated within the SM without introducing new physics. The seesaw mechanism offers an elegant explanation for these masses, naturally linking them with lepton-number-violating physics at high energy scales 111 Minkowski:1977sc , for the Minkowski’s original Type I framework, the formulations by Yanagida Yanagida:1979as and by Gell-Mann, Ramond, and Slansky GellMann:1979vx , and the left–right symmetric / Type II versions developed by Mohapatra and Senjanović Mohapatra:1980yp . . Moreover, the same framework can account for the observed matter–antimatter asymmetry of the universe via leptogenesis Fukugita:1986hr , providing a unified picture of neutrino mass generation and baryogenesis Addazi:2023majorsignals ; Addazi:2019footprints ; DiBari:2021majoronGW ; Madge:2018leptophilic ; Bosch:2023gauged .
@@
This work focuses on the theoretical and phenomenological implications of such a transition, particularly in the context of upcoming GW experiments such as LISA LISA:2017pwj , DECIGO Kawamura:2011zz , and BBO Crowder:2005nr . A detection of the predicted signal would provide direct evidence for a high-scale cosmological phase transition, indirect confirmation of lepton number violation in the neutrino sector, and strong support for the seesaw mechanism as the origin of neutrino masses. Even in the absence of a signal, the resulting constraints would have significant implications for the viable parameter space of lepton-number-breaking models.
@@ 2.1 $U(1)$ case
The vacuum expectation value (vev) $v=\langle\phi\rangle$, the scalar mass $m_{\phi}$, and the vacuum energy $V_{0}$ relate to $m_{0}$, $m_{X}$, and $\lambda_{n}$ as
@@
where these relations ensure the potential is minimized at $\phi=v$. In this work, we focus on the case $n=1$. Minimizing the potential for $n=1$ yields field-dependent masses for the scalar (radial mode) field and its associated Goldstone boson:
@@
In Fig. 3 we present the shape of the tree-level potential (left panel) and the thermally corrected (right panel) potential, where we can clearly see the thermal origin of the barrier for different scales, $m_{X}=10^{7},10^{8},10^{9}$ GeV, for $m_{0}=10^{4}$ GeV and $n_{f}=0$. This choice is motivated by collider constraints, which allow for new physics to emerge at scales starting from $m_{0}=10^{4}$ GeV. Given this lower bound, we explore what scale is required for a first-order phase transition (FOPT) associated with lepton symmetry breaking to occur. Our analysis shows that such a transition can begin to appear at around $10^{7}$ GeV. In Tab. 1 we present some benchmark points. For the thermally corrected potential we have used Eq. (1) with Eq. (2), using the complete numerical form of the thermal functions $J_{B}$ and $J_{F}$, Eq. (62), and the thermal energy corrections given in Appendix B.
@@ 2.4 $SU(2)\longrightarrow\mathds{1}$
$\left(D_{\mu}\langle\Phi\rangle\right)^{\dagger}D^{\mu}\langle\Phi\rangle=\frac{g^{2}v^{2}}{8}A_{\mu}^{a}A^{a\mu}=\frac{m_{g}^{2}}{2}A_{\mu}^{a}A^{a\mu},$ (33)
@@ 2.5 Scaling Behavior
In summary, while thermal and quantum corrections generate quartic terms that diminish with increasing $m_{X}$ , the classical structure of the potential, especially the tree-level sextic operator, drives the formation and growth of the barrier. The key effect is that the minimum shifts to larger field values as $m_{X}$ increases, amplifying the contribution of the sextic term and resulting in a barrier whose height scales as $V_{\rm Barrier}\sim m_{0}^{6}m_{X}^{4}$. This confirms that the strong first-order character of the phase transition at large $m_{X}$ is fundamentally a tree-level phenomenon.
@@ 3 Results on First Order Phase Transitions and Gravitational Waves
FOPTs emerge from the interplay between thermal fluctuations restoring symmetry and quantum corrections favoring the broken phase, with effective flat potentials leading to gravitational wave signatures. The resulting spectrum is determined by nucleation, expansion, collisions, and plasma response, and is primarily controlled by the transition strength $\alpha$, the inverse duration $\beta/H$, the nucleation temperature $T_{n}$ and the bubble wall velocity . The dynamics are described through the effective potential Eq. (1), with three main regimes: non-runaway bubbles, non-runaway bubbles in plasma, and vacuum transitions with $\alpha>1$  [^5] For an up-to-date review we consulted the latest LISA Cosmology Working Group document Caprini:2024hue .. Each case exhibits different hydrodynamical behavior and relative contributions from sound waves, bubble collisions, scalar fields, and MHD turbulence, shaping the GW spectrum. We use the most up-to-date fits available, noting that further developments in simulations may reveal additional features not yet known to the community.
@@ 3.3 Case of Vacuum Domination
Note that compared to the radiation case, in this case, since $a_{*}\gg a_{n}$, the redshift can become considerable. The vacuum domination regime is particularly relevant for our flat potential scenarios at high UV scales, where $\alpha$ can reach values of $\mathcal{O}(10^{2}-10^{3})$. In these cases, the enhanced redshift between nucleation and reheating can shift the gravitational wave peak to lower frequencies, potentially bringing signals into the optimal sensitivity bands of space-based interferometers. The temperature at which the transition takes place is given by the well known criterion of Eq.  22.124 of Maggiore:2018sht .
@@ 3.4 Parameter Space Analysis
Then in Figs. (7-9) we show $\Omega_{\rm GW}(f)h^{2}$ for some values of $\alpha$, highlighting the regions that are accessible to current experiments, LIGO LIGOScientific:2016aoc , and future experiments, LISA, DECIGO, BBO, etc.
Figure 7: FOPT signatures of $U(1)$ models with $\alpha\leq 0.5$, where the lower bounds lie within experimental sensitivity. For $m_{X}=10^{7}$ GeV the minimum sizable value is $\alpha=0.08$, and for $m_{X}=10^{9}$ GeV it is $\alpha=0.09$. For illustration we also show $m_{X}=10^{9}$ GeV with $\alpha=10$, not supported by current simulations, but indicative of the model’s possible reach.
Figure 8: FOPT signatures of the $SU(2)\rightarrow U(1)$ model. For $m_{X}=10^{9}$ GeV, $n_{f}=0$, $y=0$, we plot GW densities up to $\alpha=0.5$, which correspond to reliable simulations. For $m_{X}=10^{9}$ GeV, $n_{f}=12$, $y=0.35$, we plot the density profiles for $\alpha\leq 0.5$, which overlap with DECIGO, BBO, ET. We also plot $\alpha=15$ to illustrate possible reach to AION.
Figure 9: FOPT signature of the $SU(2)\rightarrow\mathds{1}$ model that lies in an putative observable region. Values smaller than $\alpha$ are of course possible. Also we have shown the case $\alpha=5$.
@@ 3.5 Gravitational Wave Spectral Analysis
The gravitational wave spectra shown in Figures 7-9 reveal the rich phenomenology accessible through flat potential scenarios. Each spectrum exhibits the characteristic broken power-law shape with a peak frequency determined by the Hubble scale at the transition epoch and the bubble separation scale. In Fig. 7, we have plotted the FOPT signature of the $U(1)$ models for which we get a value of $\alpha\leq 0.5$ and a lower bound on the value of $\alpha$ that lies within the experimental sensitivity of an experiment shown in the plot. For $m_{X}=10^{7}$ GeV, the minimum sizable value is $\alpha=0.08$, while for $m_{X}=10^{9}$ GeV, $\alpha=0.09$. The near-independence of this threshold on the UV scale reflects the competing effects of enhanced transition strength (which increases the signal) and modified timing relationships (which can shift the peak frequency). For illustration, we have plotted the case for $m_{X}=10^{9}$ GeV, $\alpha=10$ to illustrate the possible reach of this model, however, current simulations do not support the case of $\alpha=10$ and that line should serve only as an indication of the possible reach of the model.
@@
  • Peak frequencies typically lie in the milli-Hz range for the parameter values considered, placing them squarely within LISA’s optimal sensitivity band.
@@
For the case of the FOPT signatures of the $SU(2)\rightarrow U(1)$ model, we present the results in Fig. 8. For $m_{X}=10^{9}$ GeV, $n_{f}=0$ and $y=0$, we have plotted in shades of gray, densities corresponding to values of $\alpha$ up to $0.5$, where we can trust the simulations that produce the density profile that we use (see Appendix C). For $m_{X}=10^{9}$ GeV, $n_{f}=12$ and $y=0.35$, there is potentially a better chance to observe GW than in the previous case, since for $\alpha\leq 0.5$ there are regions that overlap DECIGO, BBO, ET AND CE. For this case, we plotted $\alpha\leq 0.5$. However, to show the reach of this model we also plot the case of $\alpha=15$, which has the potential to reach AION, ET and CE sensitivity.
@@
Note that strong phase transitions, where $\alpha>1$, have the possibility of reaching the LIGO region, but current simulations are limited to small values.

### v2

@@ Abstract
Extensions of the Standard Model typically contain “flaton fields” defined as fields with large vacuum expectation values and almost flat potentials where scalar self-coupling is small or vanishes at tree level. Such potentials have been used to drive a secondary inflationary epoch after a primary phase of inflation, in what are called thermal inflation models. Although the primordial, high-scale inflationary epoch can solve the horizon and flatness problems, it does not always resolve difficulties associated with late-time relics produced in extensions of the Standard Model. These relics typically decay too late, injecting entropy and energetic particles that spoil successful predictions like Big Bang Nucleosynthesis. It is here that thermal inflation plays a crucial role: diluting unwanted relics by many orders of magnitude without erasing the baryon asymmetry or the large-scale structure set up by the earlier phase of inflation. The preferred scale for this phenomenon is in the range $10^{6}-10^{8}$ GeV if one considers supergravity, but without it, any scale above the EW scale is valid. We investigate a typical form of these potentials and determine what are the conditions for the potentials to develop a barrier such that when the flatons settle to the true minimum, the associated Gravitational Waves can be observed, focusing on first-order phase transitions from spontaneous lepton number breaking.
CQUeST-2026-0769
@@ 1 Introduction
Among the most theoretically compelling settings for such a transition is the spontaneous breaking of lepton number symmetry, which dynamically generates Majorana masses for right-handed neutrinos. The observation of neutrino oscillations Fukuda:1998mi ; Abe:2013tm has firmly established that neutrinos have nonzero masses and mix Bilenky:1998dt ; GonzalezGarcia:2002dz ; deSalas:2018bym ; deSalas:2020pgw , a fact that cannot be accommodated within the SM without introducing new physics. The seesaw mechanism offers an elegant explanation for these masses, naturally linking them with lepton-number-violating physics at high energy scales 111 Minkowski:1977sc , for the Minkowski’s original Type I framework, the formulations by Yanagida Yanagida:1979as and by Gell-Mann, Ramond, and Slansky GellMann:1979vx , and the left–right symmetric / Type II versions developed by Mohapatra and Senjanović Mohapatra:1980yp . . Moreover, the same framework can account for the observed matter–antimatter asymmetry of the universe via leptogenesis Fukugita:1986hr , providing a unified picture of neutrino mass generation and baryogenesis Addazi:2023majorsignals ; Addazi:2019footprints ; DiBari:2021majoronGW ; Madge:2018leptophilic ; Bosch:2023gauged ; Jana:2025vyb .
@@
This work provides the first systematic study connecting flat scalar potentials to neutrino mass generation via the seesaw mechanism, with a detailed analysis of symmetry-breaking patterns in both $\mathrm{U}(1)$ and $\mathrm{SU}(2)$ frameworks. We investigate the theoretical and phenomenological implications of the associated high-scale cosmological phase transition, emphasizing its detectability in forthcoming gravitational-wave observatories such as LISA LISA:2017pwj , DECIGO Kawamura:2011zz , and BBO Crowder:2005nr . A striking outcome of the analysis is the emergence of a nontrivial scaling of the potential barrier, $V\propto m_{0}^{3}m_{X}$, which persists despite strong $1/m_{X}^{2}$ suppression of the sextic operator, revealing a subtle and unexpected enhancement mechanism. Observation of the predicted signal would constitute direct evidence for a high-scale phase transition, indirect confirmation of lepton-number violation in the neutrino sector, and strong support for the seesaw origin of neutrino masses, while null results would still impose powerful constraints on the parameter space of lepton-number-breaking models, underscoring the rich dynamics at the interface of neutrino physics and scalar field theory.
@@ 2.1 $U(1)$ case
The vacuum expectation value (vev) $v=\langle\phi\rangle$, the scalar mass $m_{\phi}$  [^4] That is $m_{\phi}^{2}(\phi)=\frac{\partial^{2}V}{\partial\phi^{2}}\bigg|_{\phi}.$, and the vacuum energy $V_{0}$ relate to $m_{0}$, $m_{X}$, and $\lambda_{n}$ as
@@
where these relations ensure the potential is minimized at $\phi=v$. In this work, we focus on the case $n=1$, because it gives the leading sextic term after higher-dimensional operators, and it is the simplest case capturing essential physics. The case $n=2$ corresponds to the power of eight, and its effects are suppressed in comparison to the sextic term effects. Minimizing the potential for $n=1$ yields field-dependent masses for the scalar (radial mode) field and its associated Goldstone boson:
@@
In Fig. 3 we present the shape of the tree-level potential (left panel) and the thermally corrected (right panel) potential, where we can clearly see the thermal origin of the barrier for different scales, $m_{X}=10^{7},10^{8},10^{9}$ GeV, for $m_{0}=10^{4}$ GeV and $n_{f}=0$. This choice is motivated by collider constraints, which allow for new physics to emerge at scales starting from $m_{0}=10^{4}$ GeV. Given this lower bound, we explore what scale is required for a first-order phase transition (FOPT) associated with lepton symmetry breaking to occur. Our analysis shows that such a transition can begin to appear at around $10^{7}$ GeV. In Tab. 1 we present some benchmark points. For the thermally corrected potential we have used Eq. (1) with Eq. (2), using the complete numerical form of the thermal functions $J_{B}$ and $J_{F}$, Eq. (64), and the thermal energy corrections given in Appendix B. We have employed our own code based on CosmoTransitions Wainwright:2011kj functions.
For the breaking of $U(1)$, a well–known consequence of Abelian symmetry breaking is the generation of cosmic strings. In our case, such strings do not lead to any observational issues since their tension is far below current constraints. The CMB bound requires $G\mu<10^{-7}$, while for the relevant range of symmetry–breaking scales considered in this work, $v\sim 10^{6}\!-\!10^{9}\,\text{GeV}$, the resulting tensions are $G\mu\sim 10^{-26}\!-\!10^{-20}$. These signals are too faint to be seen by current and proposed experiments in the frequencies that are the interest of this work (above $10^{-5}$ Hz) and see Figs. 7-9.
@@ 2.4 $SU(2)\longrightarrow\mathds{1}$
$\left(D_{\mu}\langle\Phi\rangle\right)^{\dagger}D^{\mu}\langle\Phi\rangle=\frac{g^{2}v^{2}}{8}A_{\mu}^{a}A^{a\mu}=\frac{m_{b}^{2}}{2}A_{\mu}^{a}A^{a\mu},$ (33)
@@ 2.5 Scaling Behavior
In summary, while thermal and quantum corrections generate quartic terms that diminish with increasing $m_{X}$ , the classical structure of the potential, especially the tree-level sextic operator, drives the formation and growth of the barrier. The key effect is that the minimum shifts to larger field values as $m_{X}$ increases, amplifying the contribution of the sextic term and resulting in a barrier whose height scales as $V_{\rm Barrier}\sim m_{0}^{3}m_{X}$ , Eq. (45). This confirms that the strong first-order character of the phase transition at large $m_{X}$ is fundamentally a tree-level phenomenon.
@@ 3 Results on First Order Phase Transitions and Gravitational Waves
FOPTs emerge from the interplay between thermal fluctuations restoring symmetry and quantum corrections favoring the broken phase, with effective flat potentials leading to gravitational wave signatures. The resulting spectrum is determined by nucleation, expansion, collisions, and plasma response, and is primarily controlled by the transition strength $\alpha$, the inverse duration $\beta/H$, the nucleation temperature $T_{n}$ and the bubble wall velocity . The dynamics are described through the effective potential Eq. (1), with three main regimes: non-runaway bubbles, non-runaway bubbles in plasma, and vacuum transitions with $\alpha>1$  [^6] For an up-to-date review of FOPT we consulted the latest LISA Cosmology Working Group document Caprini:2024hue . For general reviews on FOPT, see, for example, Hindmarsh:2021l ; Caprini:2020LI ; Athron:2023xlk . Each case exhibits different hydrodynamical behavior and relative contributions from sound waves, bubble collisions, scalar fields, and MHD turbulence, shaping the GW spectrum. We use the most up-to-date fits available, noting that further developments in simulations may reveal additional features not yet known to the community.
@@ 3.3 Case of Vacuum Domination
Note that compared to the radiation case, in this case, since $a_{*}\gg a_{n}$, the redshift can become considerable. The vacuum domination regime is particularly relevant for our flat potential scenarios at high UV scales, where $\alpha$ can reach values of $\mathcal{O}(10^{2}-10^{3})$. In these cases, the enhanced redshift between nucleation and reheating can shift the gravitational wave peak to lower frequencies, potentially bringing signals into the optimal sensitivity bands of space-based interferometers. The temperature at which the transition takes place is given by the well known criterion of Eq. 22.124 of Maggiore:2018sht . In the next section we compare the distinctive features of the strong phase transitions, Fig. 10, in comparison with medium and weak transitions, Fig. 7-Fig. 9.
@@ 3.4 Parameter Space Analysis
Then in Figs. (7-11) we show $\Omega_{\rm GW}(f)h^{2}$ for some values of $\alpha$, highlighting the regions that are accessible to future experiments, LISA, DECIGO, BBO, etc.
Figure 7: FOPT signatures of $U(1)$ models with $\alpha\leq 0.5$, where the lower bounds on the peak frequency, lie within experimental sensitivity. For $m_{X}=10^{7}$ GeV, the blue shaded area corresponds to $0.08\leq\alpha\leq 0.5$, and for $m_{X}=10^{9}$ GeV it is $0.09\leq\alpha\leq 0.5$.
Figure 8: FOPT signatures of the $SU(2)\rightarrow U(1)$ model. For $m_{X}=10^{9}$ GeV, $n_{f}=0$, $y=0$, we plot GW densities for $\alpha=0.23$ up to $\alpha=0.5$ (blue shaded region), which correspond to reliable simulations. For $m_{X}=10^{9}$ GeV, $n_{f}=12$, $y=0.35$, we plot the density profiles for $\alpha=0.21\leq\alpha\leq 0.5$, which overlap with DECIGO, BBO, ET.
Figure 9: FOPT signatures of the $SU(2)\rightarrow\mathds{1}$ model that lie in an putative observable regions. The shaded graded orange area corresponds to $0.3\leq\alpha\leq 0.5$.
@@ 3.5 Gravitational Wave Spectral Analysis
The gravitational wave spectra shown in Figures 7-9 reveal the rich phenomenology accessible through flat potential scenarios. Each spectrum exhibits the characteristic broken power-law shape with a peak frequency determined by the Hubble scale at the transition epoch and the bubble separation scale. In Fig. 7, we have plotted the FOPT signature of the $U(1)$ models for which we get a value of $\alpha\leq 0.5$ and a lower bound on the value of $\alpha$ that lies within the experimental sensitivity of an experiment shown in the plot. For $m_{X}=10^{7}$ GeV, the minimum sizable value is $\alpha=0.08$, while for $m_{X}=10^{9}$ GeV, $\alpha=0.09$. The near-independence of this threshold on the UV scale reflects the competing effects of enhanced transition strength (which increases the signal) and modified timing relationships (which can shift the peak frequency).
@@
For the case of the FOPT signatures of the $SU(2)\rightarrow U(1)$ model, we present the results in Fig. 8. For $m_{X}=10^{9}$ GeV, $n_{f}=0$ and $y=0$, we have plotted in shades of gray, densities corresponding to values of $\alpha$ up to $0.5$, where we can trust the simulations that produce the density profile that we use (see Appendix C). For $m_{X}=10^{9}$ GeV, $n_{f}=12$ and $y=0.35$, there is potentially a better chance to observe GW than in the previous case, since for $\alpha\leq 0.5$ there are regions that overlap DECIGO, BBO, ET and CE. For this case, we plotted $\alpha\leq 0.5$.
@@
For very strong phase transitions, $\alpha\gg 1$, dominated by by the vacuum energy, we follow Caprini:2024gyk . In this case the GW, are mainly generated by the scalar field, since bubble walls interact weakly with the plasma and most energy goes into wall motion. Historically, GW production in this regime was estimated using the envelope approximation, which neglects contributions from regions where bubble walls have collided. However, recent studies and simulations indicate that these collided wall regions play a crucial role, as walls cannot lose momentum abruptly, making previous estimates of the stochastic GW background incomplete. For completeness, we present the template that we use in Eq. (121).
In Fig. 13 we present two examples of profiles for values of $\alpha$ within the LISA observation sensitivity that correspond to this case.
Figure 10: Example of the $\alpha\gg 1$ case, for some cases of $SU(2)\rightarrow U(1)$. The first case, $\alpha=755$, corresponds to the case of $m_{X}=10^{9}$ GeV, while the second case, $\alpha=60.7$ to the case of $m_{X}=10^{10}$ GeV.
The cases plotted in Fig. 13 correspond indeed to lines 5 and 8 of Tab. 4, we will talk about it in Section (4.1). Note from Tab. 1-Tab. 2 that in each case there could be a case of vacuum domination ($\alpha\gg 1$).
Figure 11: One-dimensional plots of $\alpha$ and $\beta$ versus $g$. Left: $m_{X}=10^{7}$ GeV, $n_{f}=0$, for $U(1)$ and the fully broken $SU(2)$ ($SU(2)\rightarrow\mathds{1}$). Right: $m_{X}=10^{9}$ GeV for $n=0$, for $U(1)$ and $SU(2)\rightarrow U(1)$. A zoom highlights small $\beta/H_{*}$. $\alpha$ increases with the scale, consistent with the scaling behavior in Section 2.5, and different $SU(2)$ breaking patterns are chosen to respect BBN bounds.
Figure 12: One-dimensional plots of the corresponding peak frequency for the cases presented in Fig. 11, for $m_{X}=10^{7}$ GeV (top) and for the case of $m_{X}=10^{9}$ GeV (bottom) . For $m_{X}=10^{7}$ GeV, all the plotted range of $g$ is dominated by sound waves, while for $m_{X}=10^{9}$ GeV, for $g<0.6$ the regime corresponds to sound waves domination and $g\geq 0.6$ to bubble collisions domination.
In Fig. 11 we plot one dimensional plots showing the dependence of $\alpha$ and $\beta$ on $g$ for (left) $m_{X}=10^{7}$ GeV for $n_{f}=0$ for the $U(1)$ case and the maximally $SU(2)$ broken case, that is $SU(2)\rightarrow\mathds{1}$, and (right) $m_{X}=10^{9}$ GeV also for $n_{f}=0$. For this last case we have also plotted the case of $SU(2)\rightarrow U(1)$ (for $n_{f}=0$). Note the zoom insert that we have added such as to show the small values of $\beta/H_{*}$. For the $U(1)$ case, we can make a direct comparison in both cases. As the scale grows, the size of $\alpha$ increases significantly, as expected from the scaling behaviour we have described in Section 2.5. The reason we plot for one case $SU(2)\rightarrow\mathds{1}$ and for the other $SU(2)\rightarrow U(1)$ is because for the first case, a scale of $m_{X}=10^{9}$ GeV gives already values of $\alpha$ that saturate BBN bounds.
In Fig. 12 we plot the peak frequency for the cases of Fig. 11. For the scale $m_{X}=10^{7}$ GeV, all the range considered for $g$ lies in the weak to medium $\alpha$ regime, and therefore the overall peak frequency of the signal corresponds to the sound wave frequency, $f_{\rm sw.}$, Eq. (117). For the cases of $m_{X}=10^{9}$ GeV presented in Fig. 11, values of $g<0.6$ correspond to the sound wave regime while $g\geq 0.6$ to the bubble collisions regime, Eq. (124); for clarity, we displayed them in different figures. We did not plot $f_{\rm turb.}$ as it is a subdominant contribution in both cases, and our intention was to plot the peak frequency. We can see that the larger the gauge group, the greater the redshift (the peak frequency is lower for a fixed $g$). In these plots (Fig. 12-Fig. 11) we have plotted everything with $v_{\rm sw.}=1$ as this represents the upper limit. Other parameters entering into the computation of peak frequencies, see Appendix C, have been taken into account. There, we give the details of the values that we use for $\epsilon$ and the efficiency factor $\kappa_{\nu}$.